OpenStudy (anonymous):

lim (square root (x + square root (x + square root (x))) - square root (x)) x>+∞

OpenStudy (ikram002p):

\[\lim (\sqrt{x+\sqrt{x+\sqrt{x}}}-\sqrt{x}) \] when x apro to infinite is this the Qs ??

OpenStudy (anonymous):

yes

OpenStudy (ikram002p):

ok when its a polinomial lim f(x) = lim x^ the pigest power givin,,,

OpenStudy (anonymous):

you can solve it helps me not???

OpenStudy (ikram002p):

oh ya , just second

OpenStudy (ikram002p):

but lets do it togethor try to multibly f(x) with \[\frac{ \sqrt{x+\sqrt{x+x}}+\sqrt{x} }{ \sqrt{x+\sqrt{x+x}}+\sqrt{x} } \]

OpenStudy (ikram002p):

got it untill now ??

OpenStudy (ikram002p):

ok i'll supose that u did then the limit would be \[\lim \frac{ \sqrt{x+\sqrt{x+ \sqrt{x}}}-\sqrt{x} \times \sqrt{x+\sqrt{x+ \sqrt{x}}}+\sqrt{x} }{ \sqrt{x+\sqrt{x+ \sqrt{x}}}+\sqrt{x} }\]

OpenStudy (ikram002p):

use the concepts of "Difference between the two squares" so wat do u think r u still stuck in it ??

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

i'm so sorry..but i'm not understand....